Database last updated on Mon, 26 Jan 2026 06:40:49 +0000 by localhost
Instrument constant changes for the week before 2025324 (3 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 2 rows)
Non-critical warnings from b-files' comments for days between 2025324 and 2025324 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 323 06 14 32 Reset Warm -> Tepid from 2025 323 06 14 26 until 2025 323 06 14 26 . Counts: 1,0 | 1 | 3 | B32425.203 | CHURC | #203 | 06:10:26 | | 1 | -1 |
| rl: | 2025 324 10 53 14 New value set for watermark SUPPLY.VOLTAGE.HIGH | 2 | 3 | B32425.203 | CHURC | #203 | 06:10:29 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | ed | fz | hg | hp | rs | sl | uv | zs |
| 324 | CHURCHILL | #203 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.07599) | 2 | 3 | 33 | 13 | 12 | 2 | 8 | 20 | 38 |
(0 sec)
Relative and absolute humidity for day 2025324 (2 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 203 | 6.9 | 1.1 |